OpenStudy (anonymous):

Could someone check my work, please? These two are my original solution x=9y+3 -3x+27y=-9 Below is my work -3(9y+3)+27y=-9 -27y+9+27y=-9 0+9=-9 9=-9 My answer: There is no solution set

OpenStudy (mertsj):

-27y+9+27y=-9 Should be -9

OpenStudy (anonymous):

but wouldn't 9=-9 cancel the solution?

OpenStudy (anonymous):

Oh, ok. I see

OpenStudy (mertsj):

So now you get -9=-9 which is true but not very useful as far as finding x and y is concerned. But it gives you some valuable information and here it is: The two lines are the exact same line. So every point on one line is also on the other line and there is an infinite number of solutions. This type of system is called consistent and dependent.

OpenStudy (anonymous):

I need to subtract -9 to both sides. Then it would be -18? Is that true? -27y+9+27y=-9 -9 -9 0 doesn't equal -18

OpenStudy (mertsj):

-27y+9+27y=-9 -9 -9 Do you see where you wrote -27y+9+27y = -9?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Do you see that +9 right after the -27y?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

That positive 9 is supposed to be -9. Go way back and check your use of the distributive property.

OpenStudy (mertsj):

-3(9y+3)+27y=-9.........-3(9y+3)=-27y-9 -27y+9+27y=-9

OpenStudy (anonymous):

Oh SNAP....I didn't change my sign. Shoooooot! Oh K...I need to be careful

OpenStudy (mertsj):

Yep. It happens to all of us. Just today I was working on a calculus problem and I worked for 20 minutes trying to find a silly mistake just like that.